OpenStudy (anonymous):

I need help figuring this question out. -Find the slope of a line passing through the 2 points: (6,-2) and (1,3).

OpenStudy (whpalmer4):

The slope between two points is given by the equation \[m=\frac{(y_2-y_1)}{(x_2-x_1)}\] It doesn't matter which point you choose to be (x1, y1), only that you use the same point for x1 as you do for y1 :-)

OpenStudy (anonymous):

Not sure how to do it!

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

What's not to be sure about? You've got two points, (6, -2) and (1, 3). Pick one of them, and use its points as (x1, y1). Use the other point as (x2, y2). Plug into formula. m is the slope.

OpenStudy (anonymous):

Do i divide anything?

OpenStudy (whpalmer4):

Yes, just do as the formula says. What you are figuring out is the "rise" (amount the y value changes) over the "run" (amount the x value changes). Let's say you have points (0,0) and (6,3). That's a line that rises 3 as it goes over 6. \[m = \frac{3-0}{6-0} = 1/2\] Let's do it with the points in the other order, to show that it doesn't matter:\[m=\frac{0-3}{0-6} = \frac{-3}{-6} = 1/2\] What does matter is that you use the points in the same order. If we switch them around halfway through, we get the wrong sign: \[m=\frac{3-0}{0-6} = \frac{3}{-6} = -1/2\] Now you plug in your points in the formula and get the correct answer.

OpenStudy (hba):

You just have to plug in the point If i had the point (a,b)=(x1,y1) (c,d)=(x2,y2) then, \[m=\frac{ d-b }{ c-a }\] Now you have points (x1,y1)=6,-2) and (x2,y2)= (1,3). Use the slope formula provided by @whpalmer4

OpenStudy (hba):

@barnoldwrl Yah,Tell me what do you not understand ?

OpenStudy (whpalmer4):

@barnoldwrl You said you are still stuck. What do you mean?

OpenStudy (anonymous):

I just dont understand the first thing i do.. its just too confusing.. im stupid:(

OpenStudy (hba):

Do you understand the formula, \[Slope=m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

No

OpenStudy (whpalmer4):

Okay, do you know that the convention for a point is (x, y)? The first number is the x-value, the second number is the y-value. You have one point at (6, -2), and another at (1, 3). Let the first one be (x1, y1), so x1=6, y1 = -2. The other point will be (x2, y2), so x2 = 1, y2 = 3. Can you fill in the blanks in the formula now?

OpenStudy (anonymous):

So 3- -2 is on top and on bottom is 1- 6

OpenStudy (whpalmer4):

Yes. What does that give you for the value of m?

OpenStudy (anonymous):

so would it be 5 divided by 7? and then that is the answer?

OpenStudy (whpalmer4):

1-6 = 7? Not in my experience :-)

OpenStudy (anonymous):

that is negative 5! but what i did was, i did the problem on the top, (got an answer) then did the problem on the bottom (got the answer) then divided them by eachother.....

hero (hero):

I don't think you want to see my method.

OpenStudy (anonymous):

What do you mean :p @Hero

hero (hero):

So you want to see it?

OpenStudy (anonymous):

Lol sure

hero (hero):

(6,-2) and (1,3) plug each point into y = mx + b one at a time to get: -2 = 6m + b 3 = 1m + b isolate b in each equation: b = -2 -6m b = 3 - m set b = b -2 -6m = 3 - m -2 - 3 = 6m - m -5 = 5m -5/5 = m -1 = m